# Uncovering Weaknesses in Neural Code Generation

### v1

@@
Abstract
Code generation, the task of producing source code from prompts, has seen significant advancements with the advent of pre-trained large language models (PLMs). Despite these achievements, there lacks a comprehensive taxonomy of weaknesses about the benchmark and the generated code, which risks the community’s focus on known issues at the cost of under-explored areas.
Our systematic study aims to fill this gap by evaluating five state-of-the-art PLMs—three larger models (CodeGen2.5 (7B), CodeGeeX2 (6B), GPT-4-Turbo) and two smaller ones (UnixCoder (110M) and CodeT5-base (220M))—across three popular datasets (CoNaLa, HumanEval+, and DS-1000). We assess the quality of generated code using match-based and execution-based metrics, then conduct thematic analysis to develop a taxonomy of nine types of weaknesses.
We dissected weakness distributions in both larger and smaller models, applying an extensive methodology that encompasses model-specific as well as collective analysis (union and intersection) across models. Our research uncovers three salient findings: (1) In the CoNaLa dataset, inaccurate prompts are a notable problem, causing all large models to fail in 26.84% of cases, with even higher failure rates of 40% for smaller models; (2) Missing pivotal semantics is a pervasive issue across benchmarks, with one or more large models omitting key semantics in 65.78% of CoNaLa tasks, and similarly high occurrences in HumanEval+ (66.09%) and DS-1000 (80.51%); (3) All models struggle with proper API usage, a challenge amplified by vague or complex prompts.
Our findings aim to steer researchers towards addressing specific weaknesses/challenges in code generation. Furthermore, our annotations can offer a targeted benchmark subset for detailed analysis.
Index Terms:
@@ Abstract
I Introduction
Code generation refers to the task of automatically generating source code in a specific programming language for a given prompt. Automatic code generation has emerged as a pivotal technology in the field of software development, with the potential to reduce errors associated with manual coding and drastically improve productivity. On May 3, 2023, during the Web Summit, the GitHub CEO Thomas Dohmke demonstrated the potential of AI-assisted programming by implementing a snake game within just 15 minutes  [^1] https://rio.websummit.com/sessions/wsr23/0287e463-adf7-412b-8cf1-651de1a6128d/how-to-build-an-app-in-18-minutes/. He achieved this by just asking for and copying code generated by GitHub Copilot X, an AI-powered code generation tool. He claimed that each developer can increase their productivity by 10 times with AI. Besides, many companies have launched their own AI-powered coding assistants, such as Microsoft’s Visual Studio IntelliCode [36], Github Copilot X [17], Facebook’s Aroma [34] and so on.
In recent years, a series of impressive models have been proposed, such as sequence-based [12, 50], tree-based [56][10], and pre-trained code generation models [16][40][51][26], and they are evolving at an unprecedented rate. State-of-the-art (SOTA) code generation models, particularly those based on pre-trained large language models (PLMs), have demonstrated promising performance in generating source code for a variety of programming languages and tasks. However, despite their successes, these models still fail or perform poorly for some cases [29], and understanding their limitations is crucial for driving further advancements.
Recent studies have illuminated various challenges associated with PLMs from multiple angles. Shin et al. [45] presented a thorough evaluation of six neural code generation models, scrutinizing their performance across four benchmark datasets, and offered a detailed look at model competencies in terms of accuracy, syntax and semantic correctness, as well as the practical utility of the code produced. Antonio et al. [35] assessed GitHub Copilot’s robustness, noting that varying yet semantically similar natural language descriptions can prompt different function recommendations. Adding to this, Zan et al. [61] identified three pivotal elements for successful large language models (LLMs) in natural language to code (NL2Code) tasks: extensive model size, high-quality data, and expert fine-tuning. Furthermore, Song et al. [52] dissected code generation errors across three LLMs within the HumanEval dataset, categorizing mistakes into specific semantic and syntactic issues, such as incorrect function arguments. One similar job is done by Tambon et al. [48] on another benchmark CoderEval [57].
@@ I Introduction
In this study, we aim to uncover the weaknesses in automatic code generation by conducting an in-depth analysis of the problematic codes produced by five advanced PLMs across three widely-used datasets. In particular, we first collected five SOTA PLMs, including UnixCoder [18], CodeT5-base [51], CodeGen2.5[40], CodeGeeX2[62], and GPT-4[41], and then evaluated the generated code quality in three public datasets, including CoNaLa [55], HumanEval+ [9] and DS-1000 [24]. Through thematic analysis on the problematic cases, we identified nine weakness types from the perspectives of both benchmark and PLMs. Furthermore, we uncovered how these weaknesses are distributed among the pairs of prompts and the source code generated by different PLMs. Through this examination, we aim to underscore the predominant weakness types unique to distinct benchmarks as well as the frailties specific to various PLMs.
@@
Our findings not only serve as a foundation for further research in this area but also offer practical guidance for the development and application of more robust and reliable code generation models.
II Research Subjects
II-A Models
Generally, code-generation models can be grouped into three categories: sequence-based, tree-based and pre-trained models [26]. And the outperformance of pre-trained models, especially the recently proposed ones, have been shown by multiple research [26, 29]. Therefore, we selected five advanced code-PLMs, including UnixCoder [18], CodeT5-base [51], CodeGen2.5[40], CodeGeeX2[62] and GPT-4[41].
  • UniXCoder (110M) [18] is a multi-layer Transformer model, and it uses AST and code comment to enhance code generation. The advantage has been shown on five code-related tasks over nine datasets.
  • CodeT5-base (220M) [51] is an encoder-decoder Transformer model, in which a novel identifier-aware pre-training task is proposed to enable better identifier recovery. It has achieved SOTA results on code understanding and generation over one popular benchmark.
  • CodeGen2.5 (7B) [39] is a multi-turn program synthesis model, which progressively accepts specifications in natural language and generates the subprograms. It outperfoms CodeGen1.0-16B-mono and CodeT5-16B-mono models with only 7B on HumanEval benchmark.
  • CodeGeeX2 (6B) represents the second generation of CodeGeeX [62], a multilingual model designed specifically for code generation tasks. Despite housing only 6 billion parameters, this advanced model exhibits superior performance in code generation when tested on two popular benchmarks - HumanEval [9] and DS-1000 [24]. Remarkably, it outperforms several larger-scale, sophisticated models, including but not limited to OpenAI Codex-12B [9], LLaMA2-70B [49], and StarCoder-15B [27].
  • GPT-4-0613(GPT-4)  [^2] https://openai.com/gpt-4[41] is the state-of-the-art fourth iteration of OpenAI’s Generative Pre-trained Transformer series. It boasts an expanded parameter count, granting it an advanced capability to comprehend and process intricate instructions, thereby establishing itself as a formidable baseline for code generation applications.
@@ II-A Models
II-B Benchmark Datasets
For a comprehensive evaluation, we choose two types of datasets based on their evaluation methods: match-based evaluation (i.e., CoNaLa) and unit-test-based (i.e., HumanEval+ and DS1000). The relevant information is presented in Table I. These datasets consist of prompt-code pairs, primarily focusing on method-level granularity for the code.
  • CoNaLa [55] [1] is a collection of 2,879 pairs of how-to question and Python code fragments mined from QA forums (e.g, Stack Overflow [2]). The original team also conducted a curated processing to improve the quality of CoNaLa by incorporating variable names and function arguments into the prompts. Given that our task is to build a taxonomy of the weakness of SOTA models and high-quality datasets, we select the rewritten version.
  • HumanEval+ is an enhanced variant of the broadly utilized benchmark, HumanEval [9]. This benchmark comprises 164 programming problems and Python functions, accompanied by unit tests. In contrast to its predecessor, HumanEval+ expands the test-cases nearly 81-fold. This substantial enhancement allows for a more comprehensive evaluation of the generated code’s functional correctness. Furthermore, it significantly mitigates potential biases that might arise from single-answer evaluations.
  • DS-1000 [24] is a code generation benchmark with 1000 problems originating from 451 Stack Overflow problems. The most advantage is that it contains prompts spanning seven widely-used Python data science libraries: NumPy, Pandas, Tensorflow, PyTorch, SciPy, Scikit-learn, and Matplotlib.
TABLE I: Benchmark Datasets Information
@@ II-B Benchmark Datasets
III Case Selection and Annotation
@@
III-A Code Generation Evaluation
@@
  • CodeBLEU [43] is also a match-based metric. It not only considers the n-gram match (i.e., BLEU), but also the syntactic match via abstract syntax trees (ASTs) and code semantic match via data-flow. We integrate BLEU-4 [42] in this present work, which measures the 4-gram similarity between the generated code and the ground-truth. It is widely employed by existing approaches to evaluate the quality of code generation [26, 29].
  • Pass@K calculates the ratio of resolved prompts. For one prompt, k source code samples would be generated, and the prompt is regarded as resolved if any sample can pass the all prompt-related unit tests. In this work, we use Pass@1 because users of code-PLMs usually only need a helpful one in real practical scenario and it also has been widely used in the existing studies [26, 59].
@@
Notably, the involving three larger models, i.e., CodeGen2.5, CodeGeeX2, and GPT-4 often produce supplementary comments and functions besides the answer code. For a more accurate evaluation, we applied the respective common truncation strategies to the outputs from CodeGen2.5 and CodeGeeX2, as detailed in [39, 4]. For GPT-4’s output, since the human-machine interaction process is conducted, we manually filtered out any extraneous content for a cleaner analysis.
Results. The performance of five PLMs over the three datasets is shown in Table II. We can make two observations.
@@
TABLE II: Performance of PLMs over the Three Datasets
@@
III-B Thematic Analysis of Problematic Cases
We employ a thematic analysis process [5] to build the weakness taxonomy. We five authors form the initial group. And we recruited additional participants through fourth-year undergraduate and master CS courses at Beihang University and Harbin Institute of Technology. We conducted brief interviews and selected students with more than two years of Python development experience. In the end, we chose nine more students to join our team. These students, alongside five authors, formed an annotation group with 14 participants, with three tenured CS assistant professors and 11 CS PhD/master/undergraduate students.
@@ III-B Thematic Analysis of Problematic Cases
We then identified all problematic testing samples for each PLM to be annotated. Specifically, we selected instances in CoNaLa with an EM score of less than 1, and cases in HumanEval+ and DS-1000 where the pass@1 rate was below 1. For the two smaller models, all testing samples were deemed problematic as none produced perfect source code that either exactly matched the reference code in CoNaLa or successfully passed all given test cases in HumanEval+ and DS-1000. The number of annotated samples for each PLM across the benchmarks is detailed in Table III. For the two smaller models, the number of samples from the three benchmarks is 476, 16, and 100, respectively.
@@
TABLE III: The Distribution of Annotated Sample Sizes Across Each Benchmark for Different Models.
@@ IV Weakness taxonomy
Figure 1: The Overview of the Weakness Taxonomy.
@@
IV-A Benchmark-related weaknesses
IV-A1 Prompt-related weaknesses
I: Inaccurate description. If the prompt cannot describe the intent accurately, any PLM is almost impossible to generate the accurate source code.
@@ IV-A1 Prompt-related weaknesses
Figure 2: An Example of Diverse Responses Generated by Different Models Due to Vague Prompts.
@@
Figure 3: An Example Demonstrating How Incomplete Prompts Result in Varied Outputs Across Different Models.
@@
Figure 4: Example of a Prompt with Inconsistency.
II: Too complex prompts. Intuitively, if prompts are excessively challenging to comprehend, it becomes increasingly difficult for PLMs to generate correspondingly high-quality source code.
@@
Figure 5: Example of a Overly Complex Prompt.
III. Grammar errors. We find 7 prompts with grammar errors, mainly in CoNaLa. Grammar errors make confusion for automatic prompt comprehension, and then impede the code generation.
IV-A2 Answer-related weaknesses
IV: Biases from single answer. We discovered that 33.70% problematic cases represent false negatives in CoNaLa, for instance in Fig.6. This means that although the generated source code appears different from the golden answer, it is still correct but with a different approach, such as API usage.

### v2

@@
Abstract.
Code generation, the task of producing source code from prompts, has seen significant advancements with the advent of pre-trained large language models (PLMs). Despite these achievements, there lacks a comprehensive taxonomy of weaknesses about the benchmark and the generated code, which risks the community’s focus on known issues at the cost of under-explored areas. Our systematic study aims to fill this gap by evaluating five state-of-the-art PLMs—three larger models (CodeGen2.5 (7B), CodeGeeX2 (6B), GPT-4-Turbo) and two smaller ones (UnixCoder (110M) and CodeT5-base (220M))—across three popular datasets (CoNaLa, HumanEval+, and DS-1000). We assess the quality of generated code using match-based and execution-based metrics, then conduct thematic analysis to develop a taxonomy of nine types of weaknesses. We dissected weakness distributions and uncover three findings: (1) In the CoNaLa dataset, inaccurate prompts are a notable problem, causing all large models to fail in 26.84% of cases, with even higher failure rates of 40% for smaller models; (2) Missing pivotal semantics is a pervasive issue across benchmarks, with one or more large models omitting key semantics in 65.78% of CoNaLa, 66.09% in HumanEval+ and 80.51% in DS-1000; (3) All models struggle with proper API usage, a challenge amplified by vague or complex prompts.
@@ Abstract.
1. Introduction
Code generation refers to the task of automatically generating source code in a specific programming language for a given prompt. Automatic code generation has emerged as a pivotal technology in the field of software development, with the potential to reduce errors associated with manual coding and drastically improve productivity. On May 3, 2023, during the Web Summit, the GitHub CEO Thomas Dohmke demonstrated the potential of AI-assisted programming by implementing a snake game within just 15 minutes  [^1] https://rio.websummit.com/sessions/wsr23/0287e463-adf7-412b-8cf1-651de1a6128d/how-to-build-an-app-in-18-minutes/. He achieved this by just asking for and copying code generated by GitHub Copilot X, an AI-powered code generation tool. He claimed that each developer can increase their productivity by 10 times with AI. Besides, many companies have launched their own AI-powered coding assistants, such as Microsoft’s Visual Studio IntelliCode (Microsoft, 2023), Github Copilot X (GitHub, 2023), Facebook’s Aroma (Luan et al., 2019) and so on.
In recent years, a series of impressive models have been proposed, such as sequence-based (Dong and Lapata, 2016; Vaswani et al., 2017), tree-based (Yin and Neubig, 2018)(Chen et al., 2020), and pre-trained code generation models (Feng et al., 2020)(Nijkamp et al., 2022)(Wang et al., 2021)(Li et al., 2023b), and they are evolving at an unprecedented rate. State-of-the-art (SOTA) code generation models, particularly those based on pre-trained large language models (PLMs), have demonstrated promising performance in generating source code for a variety of programming languages and tasks. However, despite their successes, these models still fail or perform poorly for some cases (Liu et al., 2020), and understanding their limitations is crucial for driving further advancements.
Recent studies have illuminated various challenges associated with PLMs from multiple angles. Shin et al. (Shin et al., 2023) presented a thorough evaluation of six neural code generation models, scrutinizing their performance across four benchmark datasets, and offered a detailed look at model competencies in terms of accuracy, syntax and semantic correctness, as well as the practical utility of the code produced. Antonio et al. (Mastropaolo et al., 2023) assessed GitHub Copilot’s robustness, noting that varying yet semantically similar natural language descriptions can prompt different function recommendations. Adding to this, Zan et al. (Zan et al., 2023) identified three pivotal elements for successful large language models (LLMs) in natural language to code (NL2Code) tasks: extensive model size, high-quality data, and expert fine-tuning. Furthermore, Song et al. (Wang et al., 2024) dissected code generation errors across three LLMs within the HumanEval dataset, categorizing mistakes into specific semantic and syntactic issues, such as incorrect function arguments. One similar job is done by Tambon et al. (Tambon et al., 2024) on another benchmark CoderEval (Yu et al., 2024).
@@ 1. Introduction
In this study, we aim to uncover the weaknesses in automatic code generation by conducting an in-depth analysis of the problematic codes produced by five advanced PLMs across three widely-used datasets. In particular, we first collected five SOTA PLMs, including UnixCoder (Guo et al., 2022), CodeT5-base (Wang et al., 2021), CodeGen2.5(Nijkamp et al., 2022), CodeGeeX2(Zheng et al., 2023), and GPT-4(OpenAI et al., 2024), and then evaluated the generated code quality in three public datasets, including CoNaLa (Yin et al., 2018), HumanEval+ (Chen et al., 2021) and DS-1000 (Lai et al., 2022). Through thematic analysis on the problematic cases, we identified nine weakness types from the perspectives of both benchmark and PLMs. Furthermore, we uncovered how these weaknesses are distributed among the pairs of prompts and the source code generated by different PLMs. Through this examination, we aim to underscore the predominant weakness types unique to distinct benchmarks as well as the frailties specific to various PLMs.
@@
Our findings not only serve as a foundation for further research in this area but also offer practical guidance for the development and application of more robust and reliable code generation models. The generated code and our annotations are accessible via https://figshare.com/s/97fc12f62929b9b88d85.
2. Research Subjects
2.1. Models
Generally, code-generation models can be grouped into three categories: sequence-based, tree-based and pre-trained models (Li et al., 2023b). And the outperformance of pre-trained models, especially the recently proposed ones, have been shown by multiple research (Li et al., 2023b; Liu et al., 2020). Therefore, we selected five advanced code-PLMs, including UnixCoder (Guo et al., 2022), CodeT5-base (Wang et al., 2021), CodeGen2.5(Nijkamp et al., 2022), CodeGeeX2(Zheng et al., 2023) and GPT-4(OpenAI et al., 2024).
  • UniXCoder (110M) (Guo et al., 2022) is a multi-layer Transformer model, and it uses AST and code comment to enhance code generation. The advantage has been shown on five code-related tasks over nine datasets.
  • CodeT5-base (220M) (Wang et al., 2021) is an encoder-decoder Transformer model, in which a novel identifier-aware pre-training task is proposed to enable better identifier recovery. It has achieved SOTA results on code understanding and generation over one popular benchmark.
  • CodeGen2.5 (7B) (Nijkamp et al., 2023) is a multi-turn program synthesis model, which progressively accepts specifications in natural language and generates the subprograms. It outperfoms CodeGen1.0-16B-mono and CodeT5-16B-mono models with only 7B on HumanEval benchmark.
  • CodeGeeX2 (6B) represents the second generation of CodeGeeX (Zheng et al., 2023), a multilingual model designed specifically for code generation tasks. Despite housing only 6 billion parameters, this advanced model exhibits superior performance in code generation when tested on two popular benchmarks - HumanEval (Chen et al., 2021) and DS-1000 (Lai et al., 2022). Remarkably, it outperforms several larger-scale, sophisticated models, including but not limited to OpenAI Codex-12B (Chen et al., 2021), LLaMA2-70B (Touvron et al., 2023), and StarCoder-15B (Li et al., 2023a).
  • GPT-4-0613(GPT-4)  [^2] https://openai.com/gpt-4(OpenAI et al., 2024) is the state-of-the-art fourth iteration of OpenAI’s Generative Pre-trained Transformer series. It boasts an expanded parameter count, granting it an advanced capability to comprehend and process intricate instructions, thereby establishing itself as a formidable baseline for code generation applications.
@@ 2.1. Models
2.2. Benchmark Datasets
For a comprehensive evaluation, we choose two types of datasets based on their evaluation methods: match-based evaluation (i.e., CoNaLa) and unit-test-based (i.e., HumanEval+ and DS1000). The relevant information is presented in Table 1. These datasets consist of prompt-code pairs, primarily focusing on method-level granularity for the code.
  • CoNaLa (Yin et al., 2018) (Con, [n. d.]) is a collection of 2,879 pairs of how-to question and Python code fragments mined from QA forums (e.g, Stack Overflow (SO, [n. d.])). The original team also conducted a curated processing to improve the quality of CoNaLa by incorporating variable names and function arguments into the prompts. Given that our task is to build a taxonomy of the weakness of SOTA models and high-quality datasets, we select the rewritten version.
  • HumanEval+ is an enhanced variant of the broadly utilized benchmark, HumanEval (Chen et al., 2021). This benchmark comprises 164 programming problems and Python functions, accompanied by unit tests. In contrast to its predecessor, HumanEval+ expands the test-cases nearly 81-fold. This substantial enhancement allows for a more comprehensive evaluation of the generated code’s functional correctness. Furthermore, it significantly mitigates potential biases that might arise from single-answer evaluations.
  • DS-1000 (Lai et al., 2022) is a code generation benchmark with 1000 problems originating from 451 Stack Overflow problems. The most advantage is that it contains prompts spanning seven widely-used Python data science libraries: NumPy, Pandas, Tensorflow, PyTorch, SciPy, Scikit-learn, and Matplotlib.
Table 1. Benchmark Datasets Information
@@ 2.2. Benchmark Datasets
3. Case Selection and Annotation
@@
3.1. Code Generation Evaluation
@@
  • CodeBLEU (Ren et al., 2020) is also a match-based metric. It not only considers the n-gram match (i.e., BLEU), but also the syntactic match via abstract syntax trees (ASTs) and code semantic match via data-flow. We integrate BLEU-4 (Papineni et al., 2002) in this present work, which measures the 4-gram similarity between the generated code and the ground-truth. It is widely employed by existing approaches to evaluate the quality of code generation (Li et al., 2023b; Liu et al., 2020).
  • Pass@K calculates the ratio of resolved prompts. For one prompt, k source code samples would be generated, and the prompt is regarded as resolved if any sample can pass the all prompt-related unit tests. In this work, we use Pass@1 because users of code-PLMs usually only need a helpful one in real practical scenario and it also has been widely used in the existing studies (Li et al., 2023b; Zan et al., 2022b).
@@
Notably, the involving three larger models, i.e., CodeGen2.5, CodeGeeX2, and GPT-4 often produce supplementary comments and functions besides the answer code. For a more accurate evaluation, we applied the respective common truncation strategies to the outputs from CodeGen2.5 and CodeGeeX2, as detailed in (Nijkamp et al., 2023; Bavarian et al., 2022). For GPT-4’s output, since the human-machine interaction process is conducted, we manually filtered out any extraneous content for a cleaner analysis.
Results. The performance of five PLMs over the three datasets is shown in Table 2. We can make two observations.
@@
Table 2. Performance of PLMs over the Three Datasets
@@
3.2. Thematic Analysis of Problematic Cases
We employ a thematic analysis process (Braun and Clarke, 2022) to build the weakness taxonomy. We five authors form the initial group. And we recruited additional participants through fourth-year undergraduate and master CS courses at Beihang University and Harbin Institute of Technology. We conducted brief interviews and selected students with more than two years of Python development experience. In the end, we chose nine more students to join our team. These students, alongside five authors, formed an annotation group with 14 participants, with three tenured CS assistant professors and 11 CS PhD/master/undergraduate students.
@@ 3.2. Thematic Analysis of Problematic Cases
We then identified all problematic testing samples for each PLM to be annotated. Specifically, we selected instances in CoNaLa with an EM score of less than 1, and cases in HumanEval+ and DS-1000 where the pass@1 rate was below 1. For the two smaller models, all testing samples were deemed problematic as none produced perfect source code that either exactly matched the reference code in CoNaLa or successfully passed all given test cases in HumanEval+ and DS-1000. The number of annotated samples for each PLM across the benchmarks is detailed in Table 3. For the two smaller models, the number of samples from the three benchmarks is 476, 16, and 100, respectively.
@@
Table 3. The Distribution of Annotated Sample Sizes Across Each Benchmark for Different Models.
@@ 4. Weakness taxonomy
Figure 1. The Overview of the Weakness Taxonomy.
@@
4.1. Benchmark-related weaknesses
4.1.1. Prompt-related weaknesses
There are three types of weaknesses associated with the prompts.
Type I: Inaccurate description. If the prompt cannot describe the intent accurately, any PLM is almost impossible to generate the accurate source code.
@@ 4.1.1. Prompt-related weaknesses
Figure 2. An Example of Diverse Responses Generated by Different Models Due to Vague Prompts.
@@
Figure 3. An Example Demonstrating How Incomplete Prompts Result in Varied Outputs Across Different Models.
@@
Figure 4. Example of a Prompt with Inconsistency.
Type II: Too complex prompts. Intuitively, if prompts are excessively challenging to comprehend, it becomes increasingly difficult for PLMs to generate correspondingly high-quality source code.
@@
Figure 5. Example of a Overly Complex Prompt.
Type III. Grammar errors. We find 7 prompts with grammar errors, mainly in CoNaLa. Grammar errors make confusion for automatic prompt comprehension, and then impede the code generation.
4.1.2. Answer-related weaknesses
Type IV: Biases from single answer. We discovered that 33.70% problematic cases represent false negatives in CoNaLa, for instance in Fig.6. This means that although the generated source code appears different from the golden answer, it is still correct but with a different approach, such as API usage.